# Mixed Gauge-Global Symmetries, Elliptic Modes, and Black Hole Thermodynamics in Hořava-Lifshitz Gravity

### v1

@@ Abstract
Physical theories are largely dictated by their symmetries. One important conjecture in this vein is that a consistent quantum gravity theory may not possess a global symmetry, but rather only local gauge symmetries, a conjecture which has recently been extended to generalized global symmetries. Historically, this conjecture arose from a heuristic argument using black hole thermodynamics and was recently proven to hold in AdS/CFT. In Hořava-Lifshitz gravity, a presumed consistent theory of quantum gravity for which there is strong evidence for both black hole thermodynamics and a holographic construction, spacetime is endowed with a preferred dynamical spacelike foliation. The theory has a leaf reparameterization symmetry that is neither global nor local gauge, hyperbolic and elliptic equations of motion, a lack of splittability, and universal horizon black hole solutions. The reparameterization symmetry is “mixed”: it is a local symmetry in one coordinate that remains global on each leaf. More broadly it is an example of both unfree and projectable gauge symmetries. We show how the corresponding charge, which has previously not been considered thermodynamically, coupled with the other features above, resolves some outstanding tensions. Hence this mixed symmetry is not precluded by horizon thermodynamics, but is instead an integral feature. This expands the types of symmetries in quantum gravity compatible with black hole thermodynamics.
@@ 1 Introduction
One of the most important characteristics of any physical theory is the symmetries under which the theory is invariant. Requiring local gauge invariance is an integral part of the standard model of particle physics, local Lorentz symmetry dictates much of the form of special relativity, and the active diffeomorphism invariance of general relativity is a key component of holography and the AdS/CFT correspondence. Understanding what symmetries are allowed in physics is therefore an important step in any effort to construct a complete theory of quantum gravity. A key longstanding conjecture about what symmetries are allowed in quantum gravity is that there are no exact global symmetries in any complete and consistent theory of quantum gravity Banks_Constraints ; Banks ; Polchinski_Dualities ; Hsin:2020mfa ; Harlow_Main ; Harlow_PRL . This conjecture originally arose from deep considerations about black hole thermodynamics Hawking_particle and, as befitting the importance of symmetries, is intimately connected to other major conjectures about the structure of any viable quantum gravity theory including the weak gravity conjecture (c.f Harlow_WGC ), various completeness conjectures (c.f. Rudelius:2020orz ), and the distance conjecture (c.f. Cordova:2022rer ). Proofs that there are no global symmetries are also possible in more limited holographic constructions Harlow_PRL ; Harlow_Main , giving a clue that the conjecture is also relevant for holographic quantum gravity in general. As well, calculations in Euclidean quantum gravity indicate that global symmetries are not respected quantum mechanically Banks ; Strominger ; KLee_Wormhole ; ABBOTT1989687 ; COLEMAN1990387 .
In the black hole argument precluding global symmetries, which we discuss further in the next section, there is a corollary argument that symmetries must furthermore be local gauge symmetries. This is evidence, independent of any particular ultraviolet completion, that in viable quantum gravity theories the relevant symmetries are indeed local gauge symmetries Banks_Constraints . In this paper we show that there is an interesting exception to this corollary in Hořava-Lifshitz gravity, a putative complete theory of quantum gravity. Since Hořava-Lifshitz gravity possesses black holes UH and apparently a good thermodynamics of the corresponding universal horizons Mechanics ; toward_thermo ; Stefano ; StefanoVisser ; raytracing such an exception may seem counter-intuitive. However, a number of seemingly disparate features of Hořava-Lifshitz gravity conspire to make this exception possible. We briefly explain these features in the next few paragraphs to provide a conceptual picture for the reader and return to each in detail in later sections.
@@
Feature: Folations and reparameterization invariance. Hořava-Lifshitz gravity Horava_Lifshitz ; Horava_Membranes is a theory of gravity where spacetime is endowed with a dynamical spacelike foliation, the leaves of which are labelled by a scalar field $T$, the khronon. The existence of this foliation allows one to write a theory of gravity that is Lifshitz in the ultraviolet and ghost-free Horava_Lifshitz ; Horava_Membranes . The ultraviolet Lifshitz symmetry of the action then permits the theory to be renormalizable Blas , which yields a diffeomorphism invariant and seemingly complete and well-behaved theory of quantum gravity. The cost of the renormalizability of this theory is thus exact Lorentz invariance; there are tight experimental constraints multimessenger ; Kostelecky , but these can be theoretically evaded via the right hierarchies and renormalization group flow Pospelov ; Status . The structural difference between Hořava-Lifshitz gravity and general relativity is the existence of the foliation. Crucially, only the foliation is added to the theory - the actual labels of the foliation, the values of $T$, are irrelevant. This introduces a symmetry in the theory: the theory must be invariant under monotonic reparameterizations of $T$, i.e. $T\rightarrow f(T)$ is a symmetry as long as $f(T)$ is monotonic. This dictates the form of the action. The reparameterization symmetry is not a global symmetry - the corresponding global symmetry would be a shift symmetry $T\rightarrow T+T_{0}$ where $T_{0}$ is a constant. However, it is also not fully local since $f(T)$ depends only on $T$ and not on the position on any one leaf of the foliation;it is, in the parlance of Hořava-Lifshitz gravity, projectable. We will consider in this paper the reparameterization symmetry as a partially gauged global symmetry, local in $T$ but global along each leaf, and the ramifications. We denote such a symmetry as a mixed gauge-global symmetry, or mixed symmetry for short. There is a corresponding charge, but this charge does not source any fully local gauge field. Therefore, the claim that there can be no fully global symmetries remains true for Hořava-Lifshitz gravity, but the further step that only local gauge symmetries are permitted fails.
@@
Feature: Universal Horizon thermodynamics. If Hořava-Lifshitz gravity had no black holes then the above mixed symmetry and charge might cause no concerns on its face. However, Hořava-Lifshitz possesses causally trapped surfaces known as universal horizons with seemingly good thermodynamics Mechanics ; toward_thermo ; Stefano ; StefanoVisser ; raytracing ; Pacilio:2017emh ; Pacilio:2017swi ; DelPorro:2023lbv . The ultraviolet Lifshitz symmetry dictates that ultraviolet field excitations travel infinitely fast. This does not mean that causality is lost; signals still must propagate to future of each leaf. However, relativistic causality is meaningless in Hořava-Lifshitz gravity. As a result, the causal horizons are not the usual black hole horizons of general relativity - for static spacetimes the event horizon is not a Killing horizon. Instead, the dynamics of the leaves of the foliation is such that they can “bend over” in static spacetimes with a central mass. While at infinity the timelike Killing vector $\chi$ in such a spacetime can be orthogonal to the leaves, if the central mass is concentrated enough the bending of the leaves will always be sufficient so that at some finite radius inside the Killing horizon the Killing vector becomes tangential to the leaves. At this radius, since every signal propagates to the future in the foliation, a trapped surface and hence causal boundary is formed: the universal horizon. This universal horizon has a good thermodynamics and shows promise for a full holographic description with boundary Lifshitz field theories Horava_Lifshitz ; Horava_Membranes ; Iranian ; CFTno ; wedges . Key for the compatibility of mixed symmetries, their charges, and black hole thermodynamics is that all leaves outside the foliation asymptote to the universal horizon for static spacetimes and extend to spatial infinity on the other side.
@@
Feature: The presence of an elliptic mode. The requirement of a foliation coupled with the fully dynamical nature of all fields leads to an additional equation of motion for the lapse $N$, which is itself a function of $T$. Importantly, this is an elliptic equation of motion. Unlike mathematically similar constraint equations, such as the Gauss constraint, this elliptic equation is not automatically preserved by the Hamiltonian; it is not a constraint. Rather, it must be imposed as an actual equation of motion on every leaf Henneaux . The existence of an elliptic mode in a theory with black holes seems immediately problematic, as elliptic modes couple boundary data on the ends of any leaf, and if the leaf penetrated any causal boundary then one would be able to probe any central singularity by looking at the boundary data on the other side. This would violate assumed properties such as cosmic censorship. However, because all the leaves outside a universal horizon asymptote to it, the boundary data for the leaf couples points at spatial infinity not to central singularity but instead to the universal horizon data itself. In this way the elliptic mode provides a new, unique mechanism by which boundary data at a causal horizon, but not inside, can be transmitted to spatial infinity. It is in this way that the charge associated with this new symmetry can theoretically be emitted from the universal horizon despite there being no local gauge field to generate a chemical potential for different local charged, emitted particles Hook:2014mla . We note that this behavior is only possible with a) the imposition of Lifshitz symmetry (which was originally required for renormalizablility) which makes the universal horizon the causal horizon and b) foliation dynamics, which is required for holography and also generates the elliptic mode needed to pull charge off a causal boundary without a local gauge field. [^1] The elliptic mode also causes the boundary reparameterization invariance to violate the criterion of splittability, which is an assumption in proofs of no global symmetry in AdS/CFT Harlow_Main ; Harlow_PRL , allowing one to evade those proofs while still perhaps maintaining a holographic description. This aspect will be dealt with in an upcoming paper.
@@
The three features above combine into a scenario where there is a mixed symmetry which is global along leaves of the foliation yet gauged in (khronon) time, a corresponding conserved charge which sources no local gauge field, causal boundaries and black hole thermodynamics, and an elliptic mode mechanism that allows horizon data and the charge to be transmitted to infinity without gauge fields. This is a novel extension of black hole thermodynamics and hints, at least from this perspective, of more symmetry structures allowed in consistent quantum gravity theories than previously understood.  [^2] Of note here is the recent surge of interest in generalized gauge theories Gaiotto - for an introduction see Gen_Sym_Review . The shift operator maps leaves into themselves, i.e. it acts intrinsically on extended objects rather than local operators. Whether the corresponding reparameterization symmetry can be embedded into the framework of gauged generalized global symmetries is currently not known. For the purposes of this paper, when we refer to global symmetries we will not be considering generalized global symmetries, but rather the usual “0-form” global symmetries that can act on local operators. The above picture, while perhaps conceptually compelling, is a large amount of work to quantitatively check and verify all aspects. In the rest of this paper we concentrate on a simple first calculation to establish the relevance of the mixed gauge-global charge for black hole thermodynamics in Hořava-Lifshitz theory. Past work on universal horizons has ignored this charge in any first law. While the corresponding thermodynamics has worked in asymptotically flat and Lifshitz cases, it fails in asymptotically AdS when the background solution is such that normal vector to the foliations is not aligned with the timelike Killing vector at spatial infinity. We show that explicitly requiring the mixed gauge-global charge to be conserved fixes this problem - non-alignment of the normal vector corresponds to an external flux that changes the mixed gauge-global charge. Requiring a first law between equilibrium solutions for all charges therefore removes this possibility and explains why the existing derivations of thermodynamics for universal horizons failed in this case. This also establishes the relevance of such a mixed symmetry for black hole mechanics, but is only a first step in a much longer line of research.
@@
The structure of the paper is as follows. We first review the standard no global symmetries logic in Section 2. In Section 3 we introduce Hořava-Lifshitz gravity, including the basic foliation structure, the relation to Einstein-aether theory, action, and notation. Of the three necessary features, the elliptic mode is introduced in 4, reparameterization invariance as a mixed symmetry and the corresponding charge is discussed in Section 5, while the third, universal horizons and their thermodynamics, is explained in Section 6. We then show how the existing thermodynamics fails in the asymptotically AdS case in Section 7. Resolving this failure using the three features is done is Section 8, and finally we conclude with some remarks and future directions in Section 9.
2 No global symmetries in quantum gravity
It has long been conjectured that consistent theories of quantum gravity do not possess continuous global symmetries. Historically, the justification has been in the form of a heuristic argument regarding black holes that radiate via Hawking radiation Hawking_particle . More recently, the conjecture has been proven rigorously in AdS/CFT by Harlow and Ooguri in Harlow_Main ; Harlow_PRL using entanglement wedge reconstruction and splittability of boundary CFTs. In this section, we will review the heuristic justification for this conjecture and introduce the tension between the conjecture and Hořava-Lifshitz theory. The conjecture is extremely relevant for quantum gravity as it is deeply related to other conjectures, such as the weak gravity conjecture, which limits the space of consistent quantum gravity theories Harlow_Main ; Harlow_PRL .
In Hořava-Lifshitz gravity, there are black holes that radiate, and strong indications that black hole thermodynamics holds. As well, it is obvious that the theory is meant to be a consistent theory of quantum gravity. However, there are symmetries, namely the reparameterization symmetry, which are not local gauge symmetries but rather mixed, and we stress that for Hořava-Lifshitz gravity the absence of global symmetries does not automatically imply that the symmetries are local gauge, even though it is a consistent theory of quantum gravity with black hole thermodynamics. Indeed, as we shall see the existence of a mixed symmetry is actually required to resolve outstanding problems with the thermodynamics. We now turn to the basics of Hořava-Lifshitz gravity on the start of the path to see how this occurs.
@@ 5 Feature : Reparameterization invariance and conserved charge
As an aside on nomenclature, functions that are functions of only $T$ are called “projectable” in Hořava-Lifshitz gravity Horava_Membranes ; Horava_Lifshitz . Indeed, originally projectability was a key aspect of the lapse in the initial construction, although this was later shown to be too restrictive Blas . Hence one could also think of mixed symmetries as “projectable” symmetries. However, at the same time the reparameterization transformation is an example of an “unfree” gauge transformation unfree . Unfree gauge transformations are those where the gauge parameter is not a local function of $t,x$ but instead satisfies some differential equation. In the case of Hořava-Lifshitz gravity, the differential equation would simply be $\vec{\nabla}\tilde{f}(x,T)=0$, where $\vec{\nabla}$ is the spatial gradient along a leaf. This, up to a meaningless global shift, restricts $\tilde{f}(x,T)$ to simply $\tilde{f}(T)$, as we have been discussing. Finally, as mentioned previosuly, whether this symmetry can be embedded into the framework of generalized global symmetries is unknown. We choose the term “mixed” in this paper to emphasize the dual global/gauge nature of the transformation without unnecessarily pigeonholing it into any of the existing conceptual frameworks it is related to.

### v2

@@ Abstract
In Hořava-Lifshitz gravity, a putative consistent theory of quantum gravity for which there is evidence for both black hole thermodynamics and a holographic construction, spacetime is endowed with a preferred dynamical spacelike foliation. The theory has a leaf reparameterization symmetry that is neither global nor local gauge, hyperbolic and elliptic equations of motion, a lack of splittability, and universal horizon black hole solutions. The reparameterization symmetry is “mixed”: it is a local symmetry in one coordinate yet global on each leaf. More broadly it is an example of both unfree and projectable gauge symmetries. The mixed symmetry and associated charge has not yet been accounted for in calculations of universal horizon thermodynamics in Hořava-Lifshitz gravity. This has led to problems, in particular the failure of the first law in a class of asymptotically AdS solutions where the normal to the leaves of the foliation is not aligned with the time translation Killing vector at infinity. We show how the dynamics of the charge corresponding to this symmetry coupled with the other features above resolves this issue. We then briefly comment how this mixed symmetry, the corresponding charge, and the elliptic equations of motion also conspire to evade recent holographic arguments for only local gauge fields in consistent theories of quantum gravity due to the lack of splittability of the elliptic equation and associated mode.
@@ 1 Introduction
One of the most important characteristics of any physical theory is the symmetries imposed on the action. Requiring local gauge invariance is an integral part of the standard model of particle physics, local Lorentz symmetry dictates much of the form of special relativity, and the active diffeomorphism invariance of general relativity is a key component of holography and the AdS/CFT correspondence. As a consequence, symmetries also play a key role in the physics of causal horizons. The first law of black hole mechanics can be viewed as an application of Noether’s theorem, the spacetime metric is affected by the gauge field charges on the black hole, etc. In this paper we discuss the consequences for black hole thermodynamics of a “mixed” symmetry, which is neither a global nor local gauge symmetry, that exists in Hořava-Lifshitz gravity, a putative complete theory of quantum gravity. Hořava-Lifshitz gravity also possesses black holes UH and strong indications of a good thermodynamics of the corresponding universal horizons Mechanics ; toward_thermo ; Stefano ; StefanoVisser ; raytracing for flat and Lifshitz asymptotics. The mixed symmetry has not so far been explicitly examined in universal horizon thermodynamics, however. Additionally, the thermodynamics for asymptotically AdS has proven problematic, in that the first law is not consistent for all asymptotically AdS spacetimes.
The existence of a mixed symmetry and a failure of black hole thermodynamics in AdS may immediately raise concerns about the overall consistency of Hořava-Lifshitz gravity. Standard lore in quantum gravity, especially in the context of AdS/CFT, states that there are no global symmetries in a consistent quantum theory of gravity, but rather only local gauge symmetries Harlow_Main ; Harlow_PRL . Hence one might suspect that the failure of black hole thermodynamics in AdS is related to the existence of this mixed symmetry, which in turn may indicate that Hořava-Lifshitz is not actually a consistent quantum gravity theory but instead belongs to the swampland. We indeed find that the mixed symmetry and the prior failures of black hole thermodynamics in AdS are intimately related. However, instead of being an indication of inconsistency, the mixed symmetry, associated charge, and other seemingly unrelated features of Hořava-Lifshitz gravity all conspire to fix the current approaches to black hole thermodynamics when properly accounted for. While the focus of this paper is on black hole thermodynamics, a side conclusion is therefore that in some quantum gravity theories mixed symmetries may actually play a key role instead of signaling problems.
We now briefly explain the three key features of Hořava-Lifshitz gravity that make this possible to provide a conceptual picture for the reader and return to each in detail in later sections.
@@
Feature: Foliations and reparameterization invariance. Hořava-Lifshitz gravity Horava_Lifshitz ; Horava_Membranes is a theory of gravity where spacetime is endowed with a dynamical spacelike foliation, the leaves of which are labelled by a scalar field $T$, the khronon. The existence of this foliation allows one to write a theory of gravity that is Lifshitz in the ultraviolet and ghost-free Horava_Lifshitz ; Horava_Membranes . The ultraviolet Lifshitz symmetry of the action then permits the theory to be renormalizable Blas , which yields a diffeomorphism invariant and seemingly complete and well-behaved theory of quantum gravity. The cost of the renormalizability of this theory is thus exact Lorentz invariance; there are tight experimental constraints multimessenger ; Kostelecky , but these can be theoretically evaded via the right hierarchies and renormalization group flow Pospelov ; Status . The structural difference between Hořava-Lifshitz gravity and general relativity is the existence of the foliation. Crucially, only the foliation is added to the theory - the actual labels of the foliation, the values of $T$, are irrelevant. This introduces a symmetry in the theory: the theory must be invariant under monotonic reparameterizations of $T$, i.e. $T\rightarrow f(T)$ is a symmetry as long as $f(T)$ is monotonic. This heavily restricts the form of the action. The reparameterization symmetry is not a global symmetry - the corresponding global symmetry would be a shift symmetry $T\rightarrow T+T_{0}$ where $T_{0}$ is a constant. However, it is also not fully local since $f(T)$ depends only on $T$ and not on the position on any one leaf of the foliation; it is, in the parlance of Hořava-Lifshitz gravity, projectable. We will consider in this paper the reparameterization symmetry as a partially gauged global symmetry, local in $T$ but global along each leaf, and the ramifications. We denote such a symmetry as a mixed gauge-global symmetry, or mixed symmetry for short. There is a corresponding charge, but this charge does not source any fully local gauge field.
@@
Feature: Universal Horizon thermodynamics Hořava-Lifshitz possesses causally trapped surfaces known as universal horizons with seemingly good thermodynamics Mechanics ; toward_thermo ; Stefano ; StefanoVisser ; raytracing ; Pacilio:2017emh ; Pacilio:2017swi ; DelPorro:2023lbv . The ultraviolet Lifshitz symmetry dictates that ultraviolet field excitations travel infinitely fast. This does not mean that causality is lost; signals still must propagate to future of each leaf. However, relativistic causality is meaningless in Hořava-Lifshitz gravity. As a result, the causal horizons are not the usual black hole horizons of general relativity - for static spacetimes the event horizon is not a Killing horizon. Instead, the dynamics of the leaves of the foliation is such that they can “bend over” in static spacetimes with a central mass. While at infinity the timelike Killing vector $\chi$ in such a spacetime can be orthogonal to the leaves, if the central mass is concentrated enough the bending of the leaves will always be sufficient so that at some finite radius inside the Killing horizon the Killing vector becomes tangential to the leaves. At this radius, since every signal propagates to the future in the foliation, a trapped surface and hence causal boundary is formed: the universal horizon. This universal horizon has a good thermodynamics and shows promise for a full holographic description with boundary Lifshitz field theories Horava_Lifshitz ; Horava_Membranes ; Iranian ; CFTno ; wedges . Key for the compatibility of mixed symmetries, their charges, and black hole thermodynamics is that all leaves outside the foliation asymptote to the universal horizon for static spacetimes and extend to spatial infinity on the other side.
@@
Feature: The presence of an elliptic mode The requirement of a foliation coupled with the fully dynamical nature of all fields leads to an additional equation of motion for the lapse $N$, which is itself a function of $T$. Importantly, this is an elliptic equation of motion. Unlike mathematically similar constraint equations, such as the Gauss constraint, this elliptic equation is not automatically preserved by the Hamiltonian; it is not a constraint. Rather, it must be imposed as an actual equation of motion on every leaf Henneaux . The existence of an elliptic mode in a theory with black holes seems immediately problematic, as elliptic modes couple boundary data on the ends of any leaf, and if the leaf penetrated any causal boundary then one would be able to probe any central singularity by looking at the boundary data on the other side. This would violate assumed properties such as cosmic censorship. However, because all the leaves outside a universal horizon asymptote to it, the boundary data for the leaf couples points at spatial infinity not to central singularity but instead to the universal horizon data itself. In this way the elliptic mode provides a new, unique mechanism by which boundary data at a causal horizon, but not inside, can be transmitted to spatial infinity. It is in this way that the charge associated with this new symmetry can theoretically be emitted from the universal horizon despite there being no local gauge field to generate a chemical potential for different local charged, emitted particles Hook:2014mla . We note that this behavior is only possible with a) the imposition of Lifshitz symmetry (which was originally required for renormalizablility) which makes the universal horizon the causal horizon and b) foliation dynamics, which is required for holography and also generates the elliptic mode needed to pull charge off a causal boundary without a local gauge field. Specific to holographic constructions, the mixed symmetry is possible because the elliptic mode violates the criterion of splittability Harlow_Main ; Harlow_PRL used in current proofs.
@@
The three features above combine into a scenario where there is a mixed symmetry which is global along leaves of the foliation yet gauged in (khronon) time, a corresponding conserved charge which sources no local gauge field, causal boundaries and black hole thermodynamics, and an elliptic mode mechanism that allows horizon data and the charge to be transmitted to infinity without gauge fields. This is a novel extension of black hole thermodynamics with qualitatively new symmetry constructions. [^1] Of note here is the recent surge of interest in generalized gauge theories Gaiotto ; Harlow_Main ; Heckman:2024oot - for an introduction see Gen_Sym_Review . The shift operator maps leaves into themselves, i.e. it acts intrinsically on extended objects rather than local ones. Whether the corresponding reparameterization symmetry can be embedded into the framework of gauged generalized global symmetries is currently not known. For the purposes of this paper, when we refer to global symmetries we will not be considering generalized global symmetries, but rather the usual “0-form” global symmetries that can act on local objects.
The above picture, while perhaps conceptually compelling, is a large amount of work to quantitatively check and verify all aspects. In the rest of this paper we concentrate on a simple first calculation to establish the relevance of the mixed gauge-global charge for black hole thermodynamics in Hořava-Lifshitz theory. We show that explicitly requiring the mixed gauge-global charge to be conserved fixes the prior problem with asymptotically AdS spacetimes - non-alignment of the normal vector corresponds to an external flux that changes the mixed gauge-global charge. Requiring a first law between equilibrium solutions for all charges therefore removes this possibility and explains why the existing derivations of thermodynamics for universal horizons failed in this case.
@@
The structure of the paper is as follows. In Section 2 we introduce Hořava-Lifshitz gravity, including the basic foliation structure, the relation to Einstein-aether theory, action, and notation. Of the three necessary features, the elliptic mode is introduced in 3, reparameterization invariance as a mixed symmetry and the corresponding charge is discussed in Section 4, while the third, universal horizons and their thermodynamics, is explained in Section 5. We then show how the existing thermodynamics fails in the asymptotically AdS case in Section 6. Resolving this failure using the three features is done in Section 7, and finally we conclude with some remarks and future directions regarding the usual narratives around global/gauge symmetries and splittability in Section 8.
@@ 4 Feature : Reparameterization invariance and conserved charge
As an aside on nomenclature, functions that are functions of only $T$ are called “projectable” in Hořava-Lifshitz gravity Horava_Membranes ; Horava_Lifshitz . Indeed, originally projectability was a key aspect of the lapse in the initial construction, although this was later shown to be too restrictive Blas . Hence one could also think of mixed symmetries as “projectable” symmetries. However, at the same time the reparameterization transformation is an example of an “unfree” gauge transformation unfree . Unfree gauge transformations are those where the gauge parameter is not a local function of $t,x$ but instead satisfies some differential equation. In the case of Hořava-Lifshitz gravity, the differential equation would simply be $\vec{\nabla}\tilde{f}(x,T)=0$, where $\vec{\nabla}$ is the spatial gradient along a leaf. This, up to a meaningless global shift, restricts $\tilde{f}(x,T)$ to simply $\tilde{f}(T)$, as we have been discussing. Finally, as mentioned previously, whether this symmetry can be embedded into the framework of generalized global symmetries is unknown. We choose the term “mixed” in this paper to emphasize the dual global/gauge nature of the transformation without unnecessarily pigeonholing it into any of the existing conceptual frameworks it is related to.